Circular No. 6382

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) M. Womack, Pennsylvania State University (PSU); L. M. Woodney,

University of Maryland (UM); M. C. Festou, Observatoire Midi Pyrenees; J. McMullin, Submillimeter Telescope Observatory; M. F. A’Hearn, UM; D. Suswal, PSU; and S. A. Stern, Southwest Research Institute, write: “We have made the first detection of CH_{3}OH emission in comet C/1995 O1. Methanol was observed via its rotational lines at 145 GHz at high S/N with the National Radio Astronomy Observatory 12-m telescope at Kitt Peak on Apr. 16, and at somewhat lower S/N on Mar. 19 and 20. During this period, the comet moved from a heliocentric distance of 5.0 (Mar. 20) to 4.7 AU (Apr. 16). The integrated line flux area was 0.060 ± 0.003 K km s^{−1} for the (3,0)–(2,0)A methanol line at 145.103 GHz on Apr. 16 and approximately the same on Mar. 19 and 20. A rotational temperature of 12 K is derived from the (3,0)–(2,0)E, (3,−1)–(2,−1)E, and (3,0)–(2,0)A^{+} lines, from which we derive an average production rate of Q(CH_{3}OH) = 1 × 10^{27} s^{−1}. The rotational lines were not measurably blue- or red-shifted from the ephemeris velocity. The CO J = 2–1 line was observed on Apr. 15 and Mar. 18, indicating Q(CO) ∼ 8 × 10^{27} s^{−1}. The ratio of the methanoland CO-production rates is near 0.13, which is nearly three times higher than that observed in C/1996 B2 (cf. IAUC 6345 and 6353).”

COMET C/1996 B2 (HYAKUTAKE) Naked-eye m_{1} and tail-length estimates: Apr. 12.25 UT, 2.8, 100^{o} (S. J.

O’Meara, Volcano, HI); 13.20, 2.0, 10^{o} (C. E. Spratt, Victoria, BC); 14.81, 2.6, 70^{o} (K. Sárneczky, Ráktanya, Hungary); 15.14, 2.1, 70^{o} (R. Keen, Mt. Thorodin, CO); 15.82, 2.1, 25^{o} (K. Hornoch, Lelekovice, Czech Republic); 16.82, 1.8, 18^{o} (M. V. Zanotta, San Marco Pass, Italy); 17.80, 1.9, 10^{o} (A. Baransky, Kiev, Ukraine); 18.08, 2.2, 10^{o} (G. W. Kronk, Troy, IL).

COMET 67P/CHURYUMOV-GERASIMENKO Visual m_{1} estimates: Feb. 11.85, 11.1 (J. Carvajal, Madrid, Spain,

0.21-m reflector); Mar. 8.83, 11.3 (V. Znojil, Brno, Czech Republic, 25×100 binoculars); 17.05, 12.1 (J. E. Bortle, Stormville, NY, 0.41-m reflector); Apr. 7.19, 12.5 (C. S. Morris, Lockwood Valley, CA, 0.26-m reflector); 15.82, 13.3 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1996 April 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green